OpenStudy (anonymous):

-2a+3=1a-21 What does a equal?

OpenStudy (anonymous):

a=8

OpenStudy (anonymous):

can you please show work because i need help

OpenStudy (anonymous):

yup -2a+3=a-21 +2a +2a = 3=3a-21 +21 +21 24=3a 24/3=3a/3 8=a

OpenStudy (anonymous):

thanks